OpenStudy (anonymous):

write an equation in slope-intercept form of the line through points (-10,-3) and (-1,1)

OpenStudy (anonymous):

you can get the slop from any two points. y = mx + b m = slope = rise/run 1 -(-3) = 4 = rise -1 - (-10) = 9 = run m = 4/9 y = 4/9x + b SUB ONE POINT 1 = 4/9(1) + b b = 5/9 therefore y = 4/9x + 5/9

OpenStudy (anonymous):

it is multiple choice and it doesn't have 5/9 as one of the choices. it is either -13/9 or +13/9.